OpenStudy (anonymous):

1 1/3 to 1 /5/6....ratio in lowest terms simplify type as fraction

OpenStudy (anonymous):

1 1/3 = 4/3 1 5/6 = 11/6 4/3:11/6 4/3(2)=8/6 8/6:11/6 8:11 8/11

OpenStudy (anonymous):

ty